Probability
SAT · Flashcards · Recall
SAT · FLASHCARDS
Card 1
of 15
0 got it
0 to review
Current
Got it
Still learning
Not seen
QUESTION
State the formula for conditional probability
P
(
A
∣
B
)
P(A|B)
P
(
A
∣
B
)
.
QUESTION
What is the sum of probabilities for all possible outcomes of an event?
QUESTION
State the formula for the probability of the complement of an event.
Tap to flip before you grade
ANSWER
P(A') = 1 - P(A). Complement probability equals 1 minus the original event's probability.
Swipe → got it · swipe ← still learning
Got it
Still learning
Flip Card
Previous card
Next card